OpenStudy (anonymous):

A few Yes/No questions: 1.Is the ordered pair (3, 2) a solution to the system? x – 2y = 5 2x – y = 7 2.Is the ordered pair (-2,1) a solution to the system? 2x – 3y = -7 3x + y = -5 3.Is the ordered pair (1,3) a solution to -8x + 2y > -6 i think i might already know the answers, but I'm not too sure

OpenStudy (anonymous):

1. no x-2y=5 3 - 2 (2) = -1 2. yes 2x - 3y =-7 2 (-2) - 3 (1) =-7 3. yes -8 (1) + 2 (3) >-6 -2 > -6

OpenStudy (anonymous):

my answers were right then, thank you so much :)

OpenStudy (anonymous):

hehe :) good luck